# Semiparametric KSD test: unifying score and distance-based approaches for goodness-of-fit testing

### v1

@@ 5.1 Testing normality
The results are summarized in Figure 1, where “min-KSD + KSD” and “MLE + KSD” curves correspond to our proposed test. First, when data are generated under the null hypothesis, the SKSD test and the specialized normality tests (Anderson-Darling and Lilliefors) achieve the nominal level without noticeable inflation. In contrast, the LRT-based with either the minimum-KSD or MLE estimator is substantially conservative. Second, under the three alternatives, the SKSD tests exhibit power comparable to that of the specialized normality tests. In some settings, such as the Gaussian mixture model, the SKSD tests even outperform both the Anderson-Darling and Lilliefors tests. The power of the LRT-based tests, however, are quite unstable across different setups. Third, the choice of estimation algorithm $\mathcal{A}$ on the nuisance parameter does lead to differences in the power of the SKSD tests although the power gap is not substantial.
5.2 Order detection of kernel exponential model
Let $q$ be a reference density on $\mathbb{R}^{d}$, and let $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be a reproducing kernel associated with RKHS $\mathcal{H}_{\kappa}$. The kernel exponential family model [canu2006kernel] is indexed by a function $f\in\mathcal{H}_{\kappa}^{d}$ and is given by
$p_{f}(x)\propto q(x)\exp\left(\left\langle f,\kappa(x,\cdot)\mathbf{1}_{d}\right\rangle_{\mathcal{H}_{\kappa}^{d}}\right),\quad x\in\mathbb{R}^{d},$
where we set $\kappa$ to be the standard Gaussian kernel and $q(x)$ is set to be the density of $N(0,9)$. In this case, $p_{f}$ is parameterized by the functional element $f$, and the normalizing constant is typically intractable. In practice, because $\mathcal{H}_{\kappa}^{d}$ is infinite dimensional, the element $f$ cannot be estimated directly. A common strategy is to approximate $f$ by a finite-rank representation with sufficient complexity, which yields accurate density approximations [strathmann2015gradient, sutherland2018efficient]. In this experiment, we therefore focus on goodness-of-fit testing for the finite-rank approximation of the kernel exponential family.
Experiment setup.
For the null model, we assume $f(\cdot)=\sum_{\ell=1}^{p}\theta_{\ell}\phi_{\ell}(\cdot)$, with rank $p$, coefficients $\theta_{\ell}\in\mathbb{R}$ and basis function $\phi_{\ell}$. The following two settings are considered: (1) fix $p=1$ for the null model and generate data from the model with various parameter $\theta_{2}$; (2) consider the null models with different rank $p$ and generate data from the model with fixed rank $p=5$ and various choices of $\theta_{2}$. For both settings, we generate the sample with metroplis-adjusted Langevin algorithm (MALA, [roberts2001optimal]) and set the sample size to be $n=200$. We use the minimum-KSD estimator to estimate the parameters in the density model. This has become a standard approach in the literature when dealing with the nuisance parameter estimation problem [matsubara2022robust, liu2024robustness]. We refer the reader to Appendix J.2 for the complete simulation setup.
Results.
Figure 2: Power curve of SKSD goodness-of-fit tests for kernel exponential family. The red dashed line indicates the test level.
$\theta_{2}$ $p=1$ $p=2$ $p=3$ $p=4$
@@ Results.
Table 2: SKSD test power for various values of $\theta_{2}$ and null model rank pp.
In the first setting, data are generated from the null model when $\theta_{2}=0$ and from the alternative otherwise. Hence, the rejection rate should be close to the nominal level $\alpha=0.05$ at $\theta_{2}=0$ and increase as $|\theta_{2}|$ grows. Figure 2 confirms this: the Type-I error is close to $\alpha$ at $\theta_{2}=0$, and the power increases with $|\theta_{2}|$, reaching nearly 1 once $|\theta_{2}|\geq 3$, illustrating that the SKSD test is consistent and powerful in this setting. In the second setting, as $p$ increases, the null model becomes a higher-rank approximation of the kernel exponential family, which is supposed to capture more information on $f$. Table 2 shows that the power is highest at $p=1$ and generally decreases with $p$. This suggests that SKSD test can be a powerful stepwise model selection strategy for the kernel exponential family model.
@@ 5.3 Graph structure detection in conditional Gaussian model
In this application, we consider quadratic interaction model in Gaussian conditional family [gelman1991note, arnold1999conditional], which is defined as follows:
$p_{\gamma}(x)\propto\exp\left(\sum_{1\leq i\neq j\leq d}\Sigma_{ij}(x^{(i)})^{2}(x^{(j)})^{2}+\sum_{k=1}^{d}\gamma_{k}^{(2)}(x^{(k)})^{2}+\sum_{l=1}^{d}\gamma_{\ell}^{(1)}x^{(\ell)}\right)$ (11)
where in $x\in\mathbb{R}^{d}$ with parameters $\theta=(\Sigma,\gamma^{(1)},\gamma^{(2)})$ and $\Sigma\in\mathbb{R}^{d\times d}$ and $\gamma^{(1)},\gamma^{(2)}\in\mathbb{R}^{d}$. The key property of this model is that the conditional distributions are in Gaussian family. Note that the conditional Gaussian model is substantially different from the Gaussian graphical model and indeed, using graphical Lasso will lead to poor performance in terms of the graph recovery task [lin2016estimation]. Note the normalizing constant in (11) is infeasible to compute as it involves high-dimensional intractable integrals. However, the score of the model can be easily computed and we refer to Appendix J.3 for the explicit form of the score function.
@@
Without loss of generality, we consider $\Sigma$ is a symmetric matrix, i.e., $\Sigma_{ij}=\Sigma_{ji}$ for all $i,j$, and make the following constraints on the parameter space: $\Sigma_{ij}\leq 0$, $\gamma_{i}^{(2)}<0$, which ensures the existence of the joint and conditional probability distributions. We want to know if a quadratic interaction model with a ring structure, i.e., $\Sigma_{ij}=0$ for all $|i-j|\neq 1\text{ or }d-1$, can fit the model well enough. When generating the data, we consider alternatives only introducing second-order interaction terms between non-adjacent nodes in the graph: $\Sigma_{ij}=0$ for all |ij|>2|i-j|>2 or |ij|<d2|i-j|<d-2. We are interested in two settings: (1) fix dimension $d=8$ and vary the magnitude of the second-order interaction $\varepsilon$. We are under the null model when $\varepsilon=0$ and under the alternative otherwise; (2) fix the magnitude of the second-order interaction terms $\varepsilon=0.5$ and vary the dimension $d$ to generate data from different distributions, which are all under the alternatives. The parameter $\varepsilon\geq 0$ (see Appendix J.3 for the definition of $\varepsilon$) controls the magnitude of the second-order interaction terms in both settings. Sampling from model (11) can be easily implemented by Gibbs sampler [arnold2001conditionally]. We use sample size $n=500$ throughout the experiments. We refer to Appendix J.3 for additional details on the experiments.
@@
For estimation, we use minimum-KSD estimator [barp2019minimum] and score matching estimator [hyvarinen2005estimation, lin2016estimation] to estimate the parameters of the interaction terms in the null model with 1D grid structure. Both estimators admit closed-form solutions, which are given in Appendix J.3. For testing, the SKSD test with parametric bootstrap cutoff is considered (Algorithm 1).
Results.
The experimental results for the two settings are summarized in Figure 3. In the first setting with fixed $d$, the power of the test increases with $\varepsilon$ for both the score-matching estimator and the minimum-KSD estimator. Under the null, the Type I error of both estimators is well controlled at the nominal level. In the second setting, the power increases with $d$. Intuitively, in higher dimensions there are more node pairs, so the cumulative signal for detecting second-order interactions is stronger as the number of interacting nodes grows. In both setups, when the signal is weak (small $\varepsilon$ or small $d$), the score-matching estimator yields slightly higher power. As the signal strengthens, however, the minimum-KSD estimator becomes more powerful. Since score matching has $O(n)$ time complexity whereas minimum-KSD has $O(n^{2})$, these results highlight a trade-off between statistical power and computational cost. They also suggest that the choice of estimation algorithm can substantially affect the power of the SKSD test. We investigate the mechanism behind this pattern in the next section.
@@ Results.
Unveiling the mystery between different estimators.
It turns out that the different estimators trade off estimation error and separation from the null in SKSD tests. To understand these patterns, we compare the root mean squared error (RMSE) of the estimators and the distance of their population projections onto the null model. Let $\Sigma^{\mathrm{proj}}_{\mathrm{KSD}}$ (resp. $\Sigma^{\mathrm{proj}}_{\mathrm{SM}}$) be the projection of the true interaction matrix $\Sigma$ onto the null model parameter space under the loss functions of minimum-KSD (resp. score-matching) estimation procedure. For each estimator $\mathcal{A}\in\{\mathrm{KSD},\mathrm{SM}\}$, we define
@@
where $\|\cdot\|_{\mathrm{F}}$ is the Frobenius norm. The RMSE reflects estimation accuracy of different estimators, while the projection distance measures how far the projected model is from the null under alternative. Lower RMSE and larger projection distance are expected to yield higher power. To deconfound the finite-sample error, we compute the population projections using $10^{4}$ Monte Carlo samples and the RMSE using $10^{3}$ independent repetitions under the simulation setup.
The resulting ratios $\mathrm{RMSE}(\mathrm{KSD})/\mathrm{RMSE}(\mathrm{SM})$ and $\mathrm{Proj\text{-}dist}(\mathrm{KSD})/\mathrm{Proj\text{-}dist}(\mathrm{SM})$ are also shown in Figure 3. In the first setting, as $\varepsilon$ increases, the RMSE ratio decreases while the projection-distance ratio increases, matching the trend of power curves: score matching is favored when $\varepsilon$ is small (RMSE dominates), and minimum-KSD gains an advantage for larger $\varepsilon$ (projection distance dominates). In the second setting, as $d$ increases, the RMSE ratio remains nearly constant while the projection-distance ratio increases, consistent with the faster growth in power for minimum-KSD relative to score matching. In other words, it is the projection distance, i.e., the distance between the projected model and the null model, that matters most in determining the relative performance of the two estimators.
@@ 6 Discussion
In this paper, we establish a simple yet powerful connection between score-based and distance-based GoF tests and IPMs through the lens of general exponentially tilted models. This framework generalizes the classical parametric score-based test to a nonparametric setting and allows us to reinterpret many existing IPM-based methods as nonparametric score-based tests. Motivated by the computational and statistical limitations of existing tests, we propose a new nonparametric score-based test, the SKSD test paired with parametric bootstrap calibration, which can be efficiently computed and is broadly applicable to a wide range of parametric models. We prove its asymptotic properties under both the null and (local) alternative hypotheses, and we demonstrate the flexibility and effectiveness of the proposed test through extensive experiments on various models, including settings in which only the score function is available.
There are many promising directions for future work. First, it would be interesting to extend our framework to discrete distributions, where the score function can be defined through the discrete Stein operator [yang2018goodness]. Second, the choice of bandwidth in the kernel function of the SKSD test is a challenging problem and can have a substantial impact on the power of the resulting test. It would be valuable to draw inspiration from recent work on improving the power of kernel-based tests by aggregating multiple kernels [schrab2022ksd, schrab2023mmd, chatterjee2025boosting]. Lastly, the null model can be more complex, going beyond the fixed-dimensional setting to high- or even infinite-dimensional cases [zhu2020distance, zhu2021interpoint, yan2023kernel, huang2023high, gao2023two], and the composite nature of the GoF problem further complicates the analysis. An important direction is to investigate how estimation error in high-dimensional settings manifests in the nonparametric GoF testing problem and how this, in turn, suggests principled solutions.
@@ A.1 Derivation of closed-form solution for KSD
Then, we show how to derive the closed-form solution of KSD in Eqn. (6). To see this, we recall the Riesz representation theorem [reed1972methods, Theorem II.4,] so that for an RKHS $\mathcal{H}$, there exists a positive definite kernel $K:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ such that the feature map $\psi_{x}\in\mathcal{H}$ for all $x\in\mathcal{X}$ satisfies $K(x,\cdot)=\psi_{x}(\cdot)$ and $K(x,y)=\langle\psi_{x},\psi_{y}\rangle_{\mathcal{H}}$. We additionally define the operation of Stein operator $\mathcal{A}_{p}$ on a real-value function (e.g., feature map) as $\xi_{p,x}(\cdot)\in\mathcal{H}^{d}$:
@@ A.3 Two general classes of estimators
    Special instances in this class include maximum likelihood estimator (MLE), maximum pseudo-likelihood estimate (MPLE) for models with intractable likelihood [besag1975statistical].
@@
    Special instances in this class include minimum Wasserstein estimator [bernton2019parameter], minimum MMD estimator[briol2019statistical, niu2023discrepancy, alquier2024universal] and minimum-KSD estimator [barp2019minimum, matsubara2022robust].
@@ A.4 Nonparametric characteristicity
Proposition 1 (Theorem 2.2 in [chwialkowski2016kernel]).
@@ A.5 Connection to moment test
The score statistic can be interpreted from the moment equation guaranteed by Stein identity, $\mathbb{E}_{\hat{\mathcal{L}}_{n}}[\mathcal{A}_{p_{\hat{\theta}_{n}}}g(X)]=0$ for any test function $g$ in the RKHS $\mathcal{H}^{d}$. In other words, (SKSD) is a nonparametric moment test. It turns out that the moment test has been widely studied in the goodness-of-fit (GoF) testing literature, especially in the econometrics community [white1996estimation].
@@ Appendix C Neyman orthogonal SKSD test
Let $r_{\theta}(X)=\nabla_{\theta}s_{\theta}(X)$ and $\mathbf{K}(x,y)\in\mathbb{R}^{d\times d}$ be the matrix-value kernel. The transform function $\widetilde{f}$ corresponds to a transformed kernel, which we refer as the Neyman orthogonal kernel [escanciano2024gaussian] defined as
@@
1 Obtain estimate $\hat{\theta}_{n}=\mathcal{E}(X)$;
2 Compute $T_{n,\widetilde{\mathbf{K}}}\equiv T_{n,\widetilde{\mathbf{K}}}(X,\hat{\theta}_{n})$ as in Eqns. (20) and (C);
3 for $b=1,2,\dots,B$ do
4   Generate wild bootstrap weight $\{w_{i}^{(b)}\}_{i\in[n]}\overset{\mathrm{i.i.d.}}{\sim}\text{Unif}\{-1,+1\}$;

### v2

@@ 5.2 Order detection of kernel exponential model
Figure 2: Power curve of SKSD tests for kernel exponential family. The black dashed line indicates the test level.
$\theta_{2}$ $r=1$ $r=2$ $r=3$ $r=4$
@@
Table 2: SKSD test power for various values of $\theta_{2}$ and null model rank $r$.
In the first setting, data are generated from the null model when $\theta_{2}=0$ and from the alternative otherwise. Hence, the rejection rate should be close to the nominal level $\alpha=0.05$ at $\theta_{2}=0$ and heuristically, increase as $|\theta_{2}|$ grows. Figure 2 confirms this: the Type-I error is close to $\alpha$ at $\theta_{2}=0$, and the power increases with $|\theta_{2}|$, reaching nearly 1 once $|\theta_{2}|\geq 3$, illustrating that the SKSD test is consistent and powerful in this setting. In the second setting, as $r$ increases and approaches $5$, the model class expands in complexity, forming increasingly higher-rank approximations of the kernel exponential family and thus capturing more structure in $f$. Table 2 shows that the power is highest at $r=1$ and generally decreases with $r$ given the same DGP. This suggests that SKSD test can be a powerful stepwise model selection strategy for the kernel exponential family model.
@@ 5.3 Graph structure detection in conditional Gaussian model
In this application, we consider quadratic interaction model in Gaussian conditional family [36, 6], which is defined as follows:
$p_{\theta}(x)\propto\exp\left(\sum_{1\leq i\neq j\leq d}\Sigma_{ij}(x^{(i)})^{2}(x^{(j)})^{2}+\sum_{k=1}^{d}\gamma_{k}^{(2)}(x^{(k)})^{2}+\sum_{l=1}^{d}\gamma_{\ell}^{(1)}x^{(\ell)}\right)$ (11)
where $x=(x^{(1)},\ldots,x^{(d)})^{\top}\in\mathbb{R}^{d}$ with parameters $\theta=(\Sigma,\gamma^{(1)},\gamma^{(2)})$. Here, $\Sigma\in\mathbb{R}^{d\times d}$ and $\gamma^{(1)},\gamma^{(2)}\in\mathbb{R}^{d}$. Without loss of generality, we consider $\Sigma$ is a symmetric matrix, i.e., $\Sigma_{ij}=\Sigma_{ji}$ for all $i,j$, and make the following constraints on the parameter space: $\Sigma_{ij}\leq 0$, $\gamma_{i}^{(2)}<0$ for any $i,j\in[d]$, which ensures the existence of the joint and conditional probability distributions.
The key property of this model is that the conditional distributions are in Gaussian family. It is an appropriate modeling choice when inference targets are conditional in nature and Gaussianity provides a tractable working approximation for conditional noise, even when the joint distribution of covariates and responses is complex or misspecified [36, 96]. Note that the conditional Gaussian model is substantially different from the Gaussian graphical model and indeed, using graphical Lasso will lead to poor performance in terms of the graph recovery task [62]. Note the normalizing constant in (11) is infeasible to compute as it involves an high-dimensional intractable integral. However, the score function of the model can be easily computed and we refer to Appendix J.3 for the explicit form of the score function.
@@
We want to investigate if a quadratic interaction model with a ring structure, i.e., $\Sigma_{ij}=0$ for all $|i-j|\neq 1\text{ or }d-1$, can fit the model well enough. When generating the data, we consider distribution introducing second-order interaction terms between non-adjacent nodes in the graph: $\Sigma_{ij}=0$ for all |ij|>2|i-j|>2 or |ij|<d2|i-j|<d-2. The parameter $\varepsilon\geq 0$ controls the magnitude of the second-order interaction terms in both settings (see Appendix J.3 for the definition of $\varepsilon$). Sampling from model (11) can be easily implemented by Gibbs sampler [7].
We are interested in two DGPs: (1) fix dimension $d=8$ and vary the magnitude of the second-order interaction $\varepsilon$. In this case, we are under the null model when $\varepsilon=0$ and under the alternative otherwise; (2) fix $\varepsilon=0.5$ and generate data from different distributions by varying the dimension $d$. In this case, all the data are generated under the alternatives. We use sample size $n=500$ throughout the experiments and refer to Appendix J.3 for additional details on the experiments.
@@
For estimation, we use minimum-KSD estimator [10] and score matching estimator [54, 62] to estimate the nuisance parameters. Both estimators admit closed-form solutions, which are given in Appendix J.3. For testing, the SKSD test with parametric bootstrap cutoff is considered (Algorithm 1).
Simulation results.
The simulation results for the two settings are summarized in Figure 3. In the first setup with fixed $d$, the power of the tests increase with $\varepsilon$ for both the score-matching and the minimum-KSD estimators. Under the null, the Type-I error of both estimators is well controlled at the nominal level. In the second setup, the power increases with $d$. This is because, intuitively, there are more node pairs in higher dimensions so that the cumulative signal for detecting second-order interactions is stronger. In both setups, when the signal is weak (small $\varepsilon$ or small $d$), the score-matching estimator yields higher power. As the signal strengthens, however, the minimum-KSD estimator becomes more powerful. These results suggest that the choice of estimation algorithm can substantially affect the power of the SKSD test. We investigate the mechanism behind this pattern in the next section.
@@
Unveiling the mystery of test power performances.
It turns out that different estimators for $\theta$ used in the simulation trade off (1) estimation error, and (2) separation distance from the null in SKSD tests. To further understand the patterns, we compare the root mean squared error (RMSE) of two estimators and the distance of their population projections onto the null model. Let $\Sigma^{\mathrm{proj}}_{\mathrm{KSD}}$ (resp. $\Sigma^{\mathrm{proj}}_{\mathrm{SM}}$) be the projection of the interaction matrix $\Sigma$ generating data onto the null model parameter space, under the loss functions used in minimum-KSD (resp. score-matching) estimation procedure. For each estimator $\mathcal{A}\in\{\mathrm{KSD},\mathrm{SM}\}$, we define
@@
where $\|\cdot\|_{\mathrm{F}}$ is the Frobenius norm. The RMSE reflects estimation accuracy of different estimators, while the projection distance measures how far the projected model is from the null under alternative. Intuitively speaking, lower RMSE and larger projection distance are expected to yield higher power. To deconfound the finite-sample error, we compute the population projections using $10^{4}$ Monte Carlo samples and the expectation in RMSE (defined as in (12)) using $10^{3}$ independent repetitions under the simulation setup.
The resulting ratios $\mathrm{RMSE}(\mathrm{KSD})/\mathrm{RMSE}(\mathrm{SM})$ and $\mathrm{Proj\text{-}dist}(\mathrm{KSD})/\mathrm{Proj\text{-}dist}(\mathrm{SM})$ are also shown in Figure 3. In the first simulation setup, as $\varepsilon$ increases, the RMSE ratio decreases while the projection-distance ratio increases, matching the trend of power curves: score matching is favored when $\varepsilon$ is small (RMSE dominates), and minimum-KSD gains an advantage for larger $\varepsilon$ (projection distance dominates). In the second setting, as $d$ increases, the RMSE ratio remains nearly constant while the projection-distance ratio increases, consistent with the faster growth in power for minimum-KSD relative to score matching. In other words, it is the projection distance, i.e., the distance between the projected model and the null model, that matters more in determining the relative performance of the two estimators, at least under the current simulation setup.
@@ 6 Discussion
In this paper, we establish a simple yet powerful connection between score-based and distance-based GoF tests through the lens of general exponentially tilted models. This framework generalizes the classical parametric score-based test to a nonparametric setting and allows us to reinterpret many existing IPM-based methods as nonparametric score-based tests. Motivated by the computational and statistical limitations of existing tests, we propose a new nonparametric score-based test, the SKSD test paired with parametric bootstrap calibration, which can be efficiently computed and is broadly applicable to a wide range of parametric models. We prove its asymptotic properties under both the null and (local) alternative hypotheses, and we demonstrate the flexibility and effectiveness of the proposed test through extensive experiments on various models, including challenging modeling settings in which only the score function is available.
There are many promising directions for future work. First, it would be interesting to extend our framework to discrete distributions, where the score function can be defined through the discrete Stein operator [100]. Second, the choice of bandwidth in the kernel function of the SKSD test is a challenging problem and can have a substantial impact on the power of the resulting test. It would be valuable to draw inspiration from recent work on improving the power of kernel-based tests by aggregating multiple kernels [81, 82, 20]. Lastly, the null model can be more complex, going beyond the fixed-dimensional setting to high- or even infinite-dimensional cases [105, 104, 99, 51, 35], and the composite nature of the GoF problem further complicates the analysis. An important direction is to investigate how estimation error in high-dimensional settings manifests in the nonparametric GoF testing problem and how this, in turn, suggests principled solutions.
@@ A.1 Derivation of closed-form solution for KSD
Then, we show how to derive the closed-form solution of KSD in Eqn. (6). To see this, we recall the Riesz representation theorem [79, Theorem II.4,] so that for an RKHS $\mathcal{H}$, there exists a positive definite kernel $K:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ such that the feature map $\psi_{x}\in\mathcal{H}$ for all $x\in\mathcal{X}$ satisfies $K(x,\cdot)=\psi_{x}(\cdot)$ and $K(x,y)=\langle\psi_{x},\psi_{y}\rangle_{\mathcal{H}}$. We additionally define the operation of Stein operator $\mathcal{A}_{p}$ on a real-value function (e.g., feature map) as $\xi_{p,x}(\cdot)\in\mathcal{H}^{d}$:
@@ A.3 Two general classes of estimators
    Special instances in this class include maximum likelihood estimator (MLE), maximum pseudo-likelihood estimate (MPLE) for models with intractable likelihood [13].
@@
    Special instances in this class include minimum Wasserstein estimator [12], minimum MMD estimator[16, 72, 2] and minimum-KSD estimator [10, 66].
@@ A.4 Nonparametric characteristicity
Proposition 1 (Theorem 2.2 in [23]).
@@ A.5 Connection to moment test
The score statistic can be interpreted from the moment equation guaranteed by Stein identity, $\mathbb{E}_{\hat{\mathcal{L}}_{n}}[\mathcal{A}_{p_{\hat{\theta}_{n}}}g(X)]=0$ for any test function $g$ in the RKHS $\mathcal{H}^{d}$. In other words, (SKSD) is a nonparametric moment test. It turns out that the moment test has been widely studied in the goodness-of-fit (GoF) testing literature, especially in the econometrics community [96].
A.6 Comparison to the approach by [18]
Concurrent work by [18] studies composite goodness-of-fit testing via the kernel Stein discrepancy (KSD) with an estimated nuisance parameter, and develops a bootstrap framework for parameter-dependent degenerate U-statistics. Their theory is stated under high-level joint weak and bootstrap convergence conditions for the plug-in estimator, the KSD U-statistic, and its parameter gradient (see their Assumptions 5-7). In particular, their bootstrap calibration is proved to be consistent only under the null, leaving it unclear whether the procedure attains nontrivial power against contiguous local alternatives. In contrast, we introduce a semiparametric KSD—motivated by the exponentially tilted/IPM perspective—and provide a self-contained bootstrap calibration that applies to any nuisance estimator admitting a uniform asymptotic linear expansion (UALE). Our bootstrap consistency result delivers a universal conditional limit under both the null and the alternative, thereby providing firm theoretical support for our local power analysis.
@@ Appendix C Neyman orthogonal SKSD test
Let $r_{\theta}(X)=\nabla_{\theta}s_{\theta}(X)$ and $\mathbf{K}(x,y)\in\mathbb{R}^{d\times d}$ be the matrix-value kernel. The transform function $\widetilde{f}$ corresponds to a transformed kernel, which we refer as the Neyman orthogonal kernel [28] defined as
@@
Obtain estimate $\hat{\theta}_{n}=\mathcal{E}(X)$;
Compute $T_{n,\widetilde{\mathbf{K}}}\equiv T_{n,\widetilde{\mathbf{K}}}(X,\hat{\theta}_{n})$ as in Eqns. (20) and (C);
for $b=1,2,\dots,B$ do
   Generate wild bootstrap weight $\{w_{i}^{(b)}\}_{i\in[n]}\overset{\mathrm{i.i.d.}}{\sim}\text{Unif}\{-1,+1\}$;